OpenStudy (anonymous):

What is the sum of the sequence: 1+2+3+4...+1001

OpenStudy (agreene):

501501 Think of it like this... 1+999 = 1000 2+998 = 1000 3+997 = 1000 etc. How many of those do you have? and how many other things are left? you'll arrive at the above.

OpenStudy (anonymous):

Awesome! Thank you

OpenStudy (anonymous):

\[\sum _{i=1}^{1001} i=501501 \]